OpenStudy (anonymous):

An object is launched at 19.6 m/s from a height of 58.8 m. The equation for the height (h) in terms of time (t) is given by h(t) = -4.9t2 +19.6t + 58.8. What is the object's maximum height? A. 58.8 B. 194.04 C. 78.4 D. 117.6

OpenStudy (anonymous):

-4.9t2 +19.6t + 58.8. do i divide the equation by -4.9? @zepdrix

OpenStudy (anonymous):

forgot to add then i use quadratic formula

OpenStudy (anonymous):

no, you differentiate the equation and find out h'(t) = 0, then substitute found t value into h(t)

zepdrix (zepdrix):

This isn't calculus level math.

zepdrix (zepdrix):

They probably want you to put the parabola in standard form, that way you can identify the vertex.

zepdrix (zepdrix):

Ya dividing by -4.9 is probably the right way to start :)

zepdrix (zepdrix):

\[\Large h(t) = -4.9(t^2+?t+?)\]

zepdrix (zepdrix):

So what do you get for those other two values when you factor the -4.9 out?

OpenStudy (anonymous):

ok I get \[t ^{2}-4t-12\]

OpenStudy (anonymous):

a(t^2-4t-12) like that right?

zepdrix (zepdrix):

\[\Large h(t) = -4.9(t^2-4t-12)\]Ya looks good so far :)

OpenStudy (anonymous):

ok What would be the next step?

zepdrix (zepdrix):

We need to turn the t terms into a perfect square.\[\Large h(t) = -4.9(\color{orangered}{t^2-4t}-12)\]

zepdrix (zepdrix):

So what value should we add to complete the square?

OpenStudy (anonymous):

i use \[(\frac{ b }{ 2 })^{2}\]

OpenStudy (anonymous):

right?

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

just 2 or is it 2a?

zepdrix (zepdrix):

We're going to be a little sneaky. We're going to do all of this inside of the brackets. So let's ignore the a for now.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i get 4

zepdrix (zepdrix):

so we should add 4 to complete the square?\[\Large h(t) = -4.9(\color{orangered}{t^2-4t+4}-12)\]Ya that sounds right! We have a small problem though, we can't simply add 4 willy nilly. What can we do to keep this equation balanced?

OpenStudy (anonymous):

ummm....not sure goona take a guess do we add to the other side too?

zepdrix (zepdrix):

Good, that's what we would normally do, right? We would add to BOTH sides to keep things balanced. There is another trick you should be aware of though: ~We can also keep the equation balanced by adding 4 AND subtracting 4. If we do that, we're keeping it balanced while at the same time changing the way it looks.

zepdrix (zepdrix):

\[\Large h(t) = -4.9(\color{orangered}{t^2-4t+4}-4-12)\]We're adding 4 to complete our square. We'll also subtract 4 to keep the equation balanced. I colored in orange the terms we'll need for our perfect square.

zepdrix (zepdrix):

Darn we shouldn't had taken the -4.9 out of the 58.8 earlier. It gives us an extra step now. I forgot about that. :) No big deal though.

zepdrix (zepdrix):

Confused on the add/subtract thing?

OpenStudy (anonymous):

NO makes sense

zepdrix (zepdrix):

Let's combine the -4 and -12.\[\Large h(t) = -4.9(\color{orangered}{t^2-4t+4}-16)\]From here, we really only want the orange part in the brackets. So let's multiply -16 and -4.9 to get it out of the brackets.

OpenStudy (anonymous):

ok soo would it be \[h(t)=78.4(t^{2}-4t+4)\]

zepdrix (zepdrix):

Woops, that shouldn't change our a value. It will be added on the outside.\[\Large h(t) = -4.9(\color{orangered}{t^2-4t+4})+(-4.9)(-16)\]\[\Large h(t) = -4.9(\color{orangered}{t^2-4t+4})+78.4\]

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

then ill find perf square of the orange part?

zepdrix (zepdrix):

You write it as a square* yes.

OpenStudy (anonymous):

adrn sep gtg thanks for help ill finish it later :D

zepdrix (zepdrix):

It will simplify down to:\[\Large \left(t+\frac{b}{2}\right)^2\]

zepdrix (zepdrix):

oh ok :)